OpenStudy (anonymous):

I Dont Know , Im Confused .. Please Help , Ill Medal :) What is the solution to the system of equations? x + y = 2 x - y = 4 ( ___ , ___ )

OpenStudy (anonymous):

@k_lynn I have no idea...Would you help them?

OpenStudy (anonymous):

you have to solve using the substitution method

OpenStudy (anonymous):

Divide each term in the equation by -2 -2y/-2 = 2/-2 then u simplify and cancel the common factors x=2-y y=2/-2 x=2-y y=-1 you are going to have to multiply by -1 x=2-1*-1 y=-1 when u multiply -1*-1 u get 1 x=2+1 y=-1 add the 2+1 =3 so its x=3, y=-1

OpenStudy (anonymous):

think that is correct